OpenStudy (anonymous):

If A= {1,2,3,4,5,6,7,8,9} a) how many subsets does A have b) how many proper subsets does A have?

jimthompson5910 (jim_thompson5910):

if a set has n elements, then it will have 2^n subsets it will have (2^n) - 1 proper subsets because you are excluding the set itself

OpenStudy (anonymous):

is this the answer?

jimthompson5910 (jim_thompson5910):

no, you need to replace n with the number of elements in set A

jimthompson5910 (jim_thompson5910):

then evaluate

OpenStudy (anonymous):

so replace it with the number 9?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

ok so the answer would be 2^9 - 1 = 511 ?

jimthompson5910 (jim_thompson5910):

for part b)

OpenStudy (anonymous):

so a is 511 right?

jimthompson5910 (jim_thompson5910):

no what you just calculated is for part b

jimthompson5910 (jim_thompson5910):

since there are 2^9 - 1 = 511 proper subsets

OpenStudy (anonymous):

ooh ok so what about a ... is it just 9?

jimthompson5910 (jim_thompson5910):

no it's when you evaluate 2^n n = 9

OpenStudy (anonymous):

ok so how do we get a?

jimthompson5910 (jim_thompson5910):

I just told you

jimthompson5910 (jim_thompson5910):

evaluate 2^n when n = 9

OpenStudy (anonymous):

511?

OpenStudy (anonymous):

512?*